# Dist Loss: Enhancing Regression in Few-Shot Region through Distribution Distance Constraint

### v1

@@ Abstract
Imbalanced data distributions are prevalent in real-world scenarios, posing significant challenges in both imbalanced classification and imbalanced regression tasks. They often cause deep learning models to overfit in areas of high sample density (many-shot regions) while underperforming in areas of low sample density (few-shot regions). This characteristic restricts the utility of deep learning models in various sectors, notably healthcare, where areas with few-shot data hold greater clinical relevance. While recent studies have shown the benefits of incorporating distribution information in imbalanced classification tasks, such strategies are rarely explored in imbalanced regression. In this paper, we address this issue by introducing a novel loss function, termed Dist Loss, designed to minimize the distribution distance between the model’s predictions and the target labels in a differentiable manner, effectively integrating distribution information into model training. Dist Loss enables deep learning models to regularize their output distribution during training, effectively enhancing their focus on few-shot regions. We have conducted extensive experiments across three datasets spanning computer vision and healthcare: IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR. The results demonstrate that Dist Loss effectively mitigates the negative impact of imbalanced data distribution on model performance, achieving state-of-the-art results in sparse data regions. Furthermore, Dist Loss is easy to integrate, complementing existing methods. Our code will be made publicly available following the review process.
@@ 1 Introduction
Imbalanced data distributions are prevalent in the real world, with certain target values being significantly underrepresented Buda et al. (2018); Liu et al. (2019). In regression tasks, conventional deep learning models tend to predict towards regions of high sample density (many-shot regions) during training to minimize overall error. This results in models that perform well on the majority of samples but exhibit significantly higher prediction errors on regions of low sample density (few-shot regions). This phenomenon severely limits the applicability of deep learning models in certain contexts, such as healthcare scenarios where minority samples often carry significant importance, and significant errors in these samples could lead to potential adverse events.
Taking the prediction of potassium concentration based on electrocardiogram (ECG) as an example, the model takes ECG signals as input and outputs the predicted potassium concentration derived from ECG signal features. Figure 1(a) illustrates the distribution of potassium concentrations in a real-world dataset, where the majority of samples fall within the normal range, and the minority of abnormal potassium concentrations (potassium concentration $\leq$ 3.5 mmol/L or $\geq$ 5.5 mmol/L) are mostly found in the few-shot region. Due to the imbalanced data distribution, traditional deep learning models tend to predict abnormal potassium concentration samples as normal to minimize overall error. However, since abnormal potassium concentrations significantly affect metabolism and cardiac function, they can lead to serious consequences such as arrhythmias or sudden death Ferreira et al. (2020); Crotti et al. (2020); Kim et al. (2023). Therefore, in clinical settings, the focus is often on accurately predicting the minority of abnormal potassium concentrations Galloway et al. (2019); Harmon et al. (2024), which is challenging for traditional deep learning models. Hence, how to effectively enhance model accuracy in few-shot regions under imbalanced data distributions is of significant importance and value.
@@
In imbalanced regression tasks, a significant issue is the substantial discrepancy between the model’s prediction distribution and the label distribution, as demonstrated in Figure 1(b). The orange histogram in the figure represents the ground truth, whereas the blue histogram shows the distribution of the model’s predictions. It is evident that the model’s predictions are primarily concentrated in the many-shot region, while very few predictions occur in the few-shot region. This observation highlights a critical consequence of imbalanced data distributions: the marked deviation between the model’s prediction distribution and the true label distribution. While prior research has mitigated the adverse effects on minority classes in imbalanced classification by integrating distribution information into the training process Feng et al. (2018); Zheng et al. (2020); Tian et al. (2020), such strategies are rarely explored in the context of imbalanced regression. Thus, it is crucial to explore whether significant prediction errors in few-shot regions, resulting from imbalanced data distributions, can be effectively mitigated by utilizing distribution information to align the prediction distribution of the model with the label distribution.
@@
Based on this concept, we introduce a novel loss function named Dist Loss, which aims to minimize the distance between model’s prediction distribution and the label distribution. Dist Loss is implemented in three key steps: (1) Generating pseudo-labels: we use kernel density estimation (KDE) Parzen (1962) to model the probability distribution of labels and generate pseudo-labels from this distribution; (2) Creating pseudo-predictions: we sort the model’s predictions to create pseudo-predictions that reflect the prediction distribution; (3) Distance approximation: we approximate the distance between the prediction and label distributions by measuring the distance between the pseudo-labels and pseudo-predictions. By optimizing both the distribution distance and sample-level prediction errors during training, Dist Loss reduces errors in individual predictions and aligns the model’s predictions with the label distribution. This approach effectively solves the distribution discrepancy introduced by imbalanced data distributions, as shown in 1(c), thus improving the accuracy in predicting few-shot regions.
@@
To validate the effectiveness of Dist Loss, we have conducted comprehensive experiments on three datasets across computer vision and healthcare: IMDB-WIKI-DIR, AgeDB-DIR, and our meticulously crafted ECG-Ka-DIR dataset. The findings indicate that our approach achieves a substantial increase in accuracy for rare samples, thereby attaining state-of-the-art (SOTA) performance. Moreover, our experiments reveal that Dist Loss can be integrated with existing techniques, culminating in further enhanced outcomes.
@@
  • We reexamine the impact of imbalanced data distributions in regression tasks from the perspective of distribution discrepancy and introduce the concept of aligning a model’s prediction distribution with the label distribution by leveraging distribution priors.
  • We propose a novel, differentiable approach for measuring the distribution distance in regression tasks, extending distribution distance optimization techniques from classification tasks to regression domains.
  • Through extensive experiments on multiple datasets, we validate the effectiveness of Dist Loss in deep imbalanced regression, achieving SOTA performance in few-shot regions.
@@ 2.2 Imbalanced regression
Unlike imbalanced classification, regression tasks can have labels that are infinite and boundless, which prevents methods designed for imbalanced classification from being directly transferred to imbalanced regression. Consequently, existing methods focus on leveraging the continuity of the label space. At the input level, methods addressing imbalanced regression primarily focus on resampling the training dataset. SMOTE Chawla et al. (2002); Torgo et al. (2013) and its variant SMOGN Branco et al. (2017) generate new samples by leveraging the differences between minority samples and their nearest neighbors. Branco et al. (2018) integrates a bagging-based ensemble method with SMOTE to mitigate the impact of imbalanced data distributions on the model. At the feature level, Yang et al. (2021) proposes feature distribution smoothing (FDS) by transferring feature statistics between nearby target bins to smooth the feature space. VIR Wang & Wang (2024) borrows data with similar regression labels to compute the variational distribution of the latent representation. Ranksim Gong et al. (2022) uses contrastive learning to bring the feature space of samples with similar labels closer and push the feature space of samples with dissimilar labels apart. ConR Keramati et al. (2024) designs positive and negative sample pairs based on label similarity, transferring the label space relationships to the feature space in a contrastive manner, too. At the model output and label level, regressor retraining (RRT) Yang et al. (2021) decouples the training of the encoder and regressor, retraining the regressor with inverse reweighting after normal encoder training. DenseLoss Steininger et al. (2021) and label distribution smoothing (LDS) Yang et al. (2021) measure label rarity through KDE and assign weights to each sample, with rare samples being assigned higher weights to enhance the model’s focus on minority samples. Balanced MSE Ren et al. (2022) leverages the training label distribution prior to restore a balanced prediction.
However, existing research on deep imbalanced regression often overlooks the significant distribution discrepancy between model’s predictions and labels, and distribution information, which has been proven effective in imbalanced classification, is rarely utilized. In contrast, our approach focuses on concurrently aligning the prediction distribution with the label distribution and reducing sample-level prediction errors during the training process. This significantly improves the model’s accuracy on few-shot regions without incurring additional computational costs or requiring meticulous hyperparameter tuning. Extensive experiments demonstrate the superiority of our approach in handling critical and informative rare samples in few-shot regions, achieving SOTA results.
@@ 3 Method
Figure 2: The presence of imbalanced data distributions introduces a noticeable discrepancy between the model’s prediction and label distributions. Dist Loss mitigates the impact of this imbalance by minimizing this discrepancy. Initially, KDE is utilized to ascertain the distribution of labels and to calculate the expected frequencies of each label within a batch, thereby generating pseudo-labels imbued with the label distribution information. For example, given the labels [1, 3, 4, 6] and their calculated expected frequencies [1, 2, 3, 1], the resulting pseudo-labels would be [1, 3, 3, 4, 4, 4, 6], where each label is repeated according to its calculated frequency. Subsequently, the model’s predictions within a batch are sorted to yield an ordered sequence of predictive values. Assuming the model’s initial predictions are [5, 2, 6, 3, 2, 7, 1], the sorted sequence, which encapsulates the prediction distribution information, would be [1, 2, 2, 3, 5, 6, 7]. Measuring the distance between these pseudo-labels and pseudo-predictions, both equipped with respective distribution information, provides an approximation of the distribution distance. Thereafter, by simultaneously optimizing the distribution distance and sample-level prediction errors during the training process, the model can effectively alleviate the negative effects of imbalanced data, significantly enhancing its accuracy in few-shot regions.
@@ 3.1 Problem setting
Let $\mathcal{D}$ be a training dataset comprising $N$ samples, denoted as $\mathcal{D}=\{(\mathbf{x}_{(i)},y_{(i)})\}_{i=1}^{N}$, where $\mathbf{x}_{(i)}\in\mathbb{R}^{d}$ represents the input and $y_{(i)}\in\mathbb{R}$ denotes the corresponding label. To facilitate processing, the continuous label space $\mathcal{Y}$ is discretized into $B$ bins of equal width: $\mathcal{Y}=\bigcup_{b=1}^{B}[y_{b},y_{b+1}),$ where $y_{b}$ is the lower bound of bin $b$, and $y_{1}<y_{2}<\cdots<y_{B}$. In subsequent discussions, for convenience, the lower bound $y_{b}$ of the bin $[y_{b},y_{b+1})$ will represent any label value $y_{(i)}$ that falls within that bin. Similarly, the model prediction space $\hat{\mathcal{Y}}$ is partitioned into bins of equal width. In practical scenarios, the width of each bin, denoted $\Delta_{y}$, indicates the minimum resolution of interest when processing the label space. For instance, in age estimation, one might set the bin width to 1, resulting in: $\Delta_{y}=y_{b+1}-y_{b}=1,\quad\forall b\in\mathcal{B}.$ Additionally, we define the probability of observing a label $y_{i}$ as $p_{i}$, and the probability distribution of labels can be estimated using KDE.
@@ 3.2 Dist Loss
One of the optimization objectives of Dist Loss is to minimize the distance between the prediction and label distributions in regression tasks. The core challenge lies in measuring the distance between these two distributions in a differentiable manner. Traditional metrics for measuring distribution distance, such as Kullback-Leibler divergence and Jensen-Shannon divergence, cannot be implemented in a differentiable form for regression tasks. Therefore, we have devised an alternative approach in the implementation of Dist Loss to realize a differentiable distribution distance measurement in regression scenarios. Specifically, we approximate the distance between the label and prediction distributions by sampling from these distributions and quantifying the differences between the sampled values to estimate the distance.
@@ 3.2.1 Calculation of Dist Loss
To generate pseudo-labels that contain label distribution information, we first randomly sample $M$ points from the label distribution. The expected frequencies of the label $y_{i}$ can be estimated by multiplying the number of sampling points $M$ by the probability of that label $p_{i}$. Based on this, we construct a sequence $\mathcal{N}_{L}=(n_{1},n_{2},\cdots,n_{B})$ to represent these expected frequencies, where $n_{i}=M\cdot p_{i}$. Each element in the obtained $\mathcal{N}_{L}$ represents the expected frequencies of the corresponding label. Since these frequencies may be fractional, we need to convert them to integers while ensuring that the sum after conversion still equals $M$. Here, we denote the converted integer sequence by $\mathcal{N}_{L^{\prime}}=(n^{\prime}_{1},n^{\prime}_{2},\cdots,n^{\prime}_{B})$. To acquire $\mathcal{N}_{L^{\prime}}$, we first take the floor of each element in $\mathcal{N}_{L}$ to obtain the sequence $\mathcal{N}_{L_{f}}=(\lfloor n_{1}\rfloor,\lfloor n_{2}\rfloor,\cdots,\lfloor n_{B}\rfloor)$. Then we calculate the difference $a$, which represents the difference between the sum of the original expected frequencies ($M$) and the sum after applying the floor function, following $a=M-\sum_{i=1}^{B}\lfloor n_{i}\rfloor$. Using the difference $a$, we construct an auxiliary sequence $\mathcal{A}$, which determines how to evenly distribute the difference to the elements of $\mathcal{N}_{L_{f}}$ to ensure the sum is $M$:
@@
Each $n^{\prime}_{i}$ is determined by adding $a_{i}$ to the corresponding element in $\mathcal{N}_{L_{f}}$, where $n^{\prime}_{i}=\lfloor n_{i}\rfloor+a_{i}$, and i $\in\mathcal{B}$. Finally, we generate the corresponding pseudo-labels $\mathcal{S}_{L}$ based on the expected frequencies, where each element $S_{L_{j}}$ is represented as:
$S_{L_{j}}=\min_{i\in\mathcal{B}}\left(y_{i}\cdot\theta\left(\sum_{k=1}^{i}n^{\prime}_{k}-j\right)\right),$
Here, $\theta(x)$ is the unit step function, which returns 1 when $x\geq 0$ and 0 otherwise. To illustrate with a specific example, assume that the label sequence is $(y_{1},y_{2},y_{3})=(4,5,6)$ and that the obtained sequence $\mathcal{N}_{L^{\prime}}$ is $(1,2,3)$. Then, the generated pseudo-labels $S_{L}$ would be $(4,5,5,6,6,6)$.
Similarly, we can perform $M$-point sampling on the prediction distribution and subsequently apply the same operations to obtain the pseudo-predictions $\mathcal{S}_{P}$, which encapsulate information about the prediction distribution.
In practice, we can consider a batch during the model training process as a random sampling event, wherein the model predictions within a batch are viewed as the sampling values of the prediction distribution. Consequently, we do not need to repeat the aforementioned process to acquire pseudo-predictions; instead, we simply sort the model predictions in the batch, which already contains the prediction information. By measuring the distance between the pseudo-predictions and the pseudo-labels, we can approximate the distance between the respective distributions. Let us denote the function $L(\cdot)$ as a measure of the distance between two sequences; then, the distribution distance can be expressed as $L(\mathcal{S}_{P},\mathcal{S}_{L})$. Furthermore, using the function $L(\cdot)$, we can simultaneously assess the sample-level prediction errors, which reflect the difference between the predicted values and the labels. Ultimately, by synchronously optimizing both the distribution distance and the sample-level prediction errors during the training process, we can mitigate the issue of distribution discrepancy, thereby addressing the challenges posed by imbalanced data distributions.

### v2

@@ Abstract
Imbalanced data distributions are prevalent in real-world scenarios, presenting significant challenges in both classification and regression tasks. This imbalance often causes deep learning models to overfit in regions with abundant data (many-shot regions) while underperforming in regions with sparse data (few-shot regions). Such characteristics limit the applicability of deep learning models across various domains, notably in healthcare, where rare cases often carry greater clinical significance. While recent studies have highlighted the benefits of incorporating distributional information in imbalanced classification tasks, similar strategies have been largely unexplored in imbalanced regression. To address this gap, we propose Dist Loss, a novel loss function that integrates distributional information into model training by jointly optimizing the distribution distance between model predictions and target labels, alongside sample-wise prediction errors. This dual-objective approach encourages the model to balance its predictions across different label regions, leading to significant improvements in accuracy in few-shot regions. We conduct extensive experiments across three datasets spanning computer vision and healthcare: IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR. The results demonstrate that Dist Loss effectively mitigates the impact of imbalanced data distributions, achieving state-of-the-art performance in few-shot regions. Furthermore, Dist Loss is easy to integrate and complements existing methods. To facilitate further research, we provide our implementation at https://github.com/Ngk03/DIR-Dist-Loss.
footnotetext: * Authors contributed equally to this research.footnotetext: Corresponding author.
@@ 1 Introduction
Imbalanced data distributions are prevalent in the real world, where certain target values are significantly underrepresented Buda et al. (2018); Liu et al. (2019). In imbalanced regression tasks, deep learning models tend to bias their predictions toward regions with abundant data (many-shot regions) to minimize overall error, resulting in significantly higher errors in sparse-data regions (few-shot regions). This phenomenon severely limits the applicability of deep learning models in certain contexts, such as healthcare scenarios where rare cases often carry significant importance, and significant errors in these samples could lead to potential adverse events.
Taking the prediction of potassium concentration from electrocardiogram (ECG) signals as an example, the model takes ECG signals as input and predicts the corresponding potassium concentration. Figure 1(a) illustrates the distribution of potassium concentrations in a real-world dataset, where the majority of samples fall within the normal range, while abnormal potassium concentrations ($\leq$ 3.5 mmol/L or $\geq$ 5.5 mmol/L) are rare and concentrated in the few-shot region. Due to this imbalance, deep learning models tend to bias their predictions towards the normal range in order to minimize overall error (Figure 1(b)). However, abnormal potassium concentrations can severely affect metabolism and cardiac function, potentially leading to life-threatening arrhythmias or sudden death Ferreira et al. (2020); Crotti et al. (2020); Kim et al. (2023). Therefore, in clinical settings, accurately predicting abnormal potassium concentrations is far more critical than accurately identifying normal levels Galloway et al. (2019); Harmon et al. (2024), yet this remains a challenge for existing deep learning models. Enhancing model accuracy in the few-shot region under imbalanced data distributions continues to be both a significant challenge and an important goal Branco et al. (2017); Steininger et al. (2021).
@@
In imbalanced regression tasks, a significant challenge lies in the substantial discrepancy between the distribution of the model’s predictions and the true label distribution, as illustrated in Figure 1(b). The orange histogram represents the ground truth, while the blue histogram depicts the distribution of predictions made by the vanilla model, which refers to a model trained without techniques specifically designed to address data imbalance (using L1 loss here). It is evident that the model predominantly predicts values within the many-shot region, with very few predictions falling within the few-shot region. This discrepancy highlights a critical consequence of imbalanced data distributions: the misalignment between the model’s predictions and the target labels. While prior research has focused on addressing the adverse effects of class imbalance in classification tasks by incorporating distributional information into the training process Feng et al. (2018); Zheng et al. (2020); Tian et al. (2020), similar strategies have been largely unexplored in imbalanced regression. Therefore, a key research direction is to investigate whether leveraging distributional information can effectively reduce prediction errors in the few-shot region by aligning the distribution of the model’s predictions with the underlying label distribution.
@@
Based on this concept, we propose Dist Loss, a novel loss function comprising two key components: distribution alignment optimization, which minimizes the discrepancy between the model’s prediction distribution and the label distribution, and sample-wise prediction optimization, which ensures accurate predictions at the individual sample level. The distribution alignment optimization is achieved through three steps: (1) Generating pseudo-labels: we apply kernel density estimation (KDE) to the label set Parzen (1962) to model the label distribution and sample pseudo-labels that reflect this distribution; (2) Constructing pseudo-predictions: we sort the model’s predictions to construct pseudo-predictions that represent the prediction distribution; (3) Measuring distribution discrepancy: we approximate the discrepancy between the prediction and label distributions by computing the distance between the pseudo-labels and pseudo-predictions. By jointly optimizing distribution alignment and sample-wise prediction accuracy, Dist Loss mitigates errors in individual predictions while ensuring that the model’s prediction distribution better conforms to the label distribution. As shown in Figure 1(c), this approach effectively addresses the distribution mismatch caused by data imbalance and has been proven to improve prediction accuracy in the few-shot regions.
@@
To validate the effectiveness of Dist Loss, we conduct comprehensive experiments on three datasets spanning computer vision and healthcare: IMDB-WIKI-DIR, AgeDB-DIR, and our meticulously curated ECG-K-DIR dataset. Our results demonstrate a substantial improvement in accuracy for rare samples, leading to state-of-the-art (SOTA) performance. Furthermore, our experiments show that Dist Loss is compatible with existing techniques, yielding further performance gains when integrated.
@@
  • We analyze the impact of imbalanced data distributions in regression from a distributional perspective and introduce the concept of aligning the model’s prediction distribution with the label distribution by leveraging distributional priors.
  • We propose a differentiable approach for measuring distribution distance in regression tasks, extending distribution alignment techniques from classification to regression.
  • Extensive experiments on diverse datasets demonstrate the effectiveness of Dist Loss in imbalanced regression, achieving SOTA performance in few-shot regions.
@@ 2.2 Imbalanced regression
Unlike classification tasks, where labels are discrete and bounded, regression tasks involve continuous and unbounded labels, and the distances between labels carry meaningful semantic information. These fundamental differences prevent methods designed for imbalanced classification from being directly applied to imbalanced regression. Existing methods for addressing imbalanced regression can be categorized into three levels: input, feature, and model output. At the input level, methods primarily focus on resampling the training dataset. SMOTE Chawla et al. (2002); Torgo et al. (2013) and its variant SMOGN Branco et al. (2017) generate new samples by interpolating between minority samples and their nearest neighbors. Branco et al. (2018) further enhances this approach by integrating a bagging-based ensemble method with SMOTE to mitigate the impact of imbalanced data distributions. At the feature level, Yang et al. (2021) introduces feature distribution smoothing (FDS), which transfers feature statistics between nearby target bins to smooth the feature space. VIR Wang & Wang (2024) borrows data with similar regression labels to compute the variational distribution of the latent representation. Ranksim Gong et al. (2022) employs contrastive learning to bring the feature representations of samples with similar labels closer while pushing apart those with dissimilar labels. Similarly, ConR Keramati et al. (2024) designs positive and negative sample pairs based on label similarity, transferring label space relationships to the feature space in a contrastive manner. At the model output and label level, regressor retraining (RRT) Yang et al. (2021) decouples the training of the encoder and regressor, retraining the regressor with inverse reweighting after normal encoder training. DenseLoss Steininger et al. (2021) and label distribution smoothing (LDS) Yang et al. (2021) measure label rarity through KDE, assigning higher weights to rare samples to enhance the model’s focus on minority samples. Balanced MSE Ren et al. (2022) leverages the training label distribution prior to restore balanced predictions.
However, existing research on imbalanced regression often overlooks the significant distribution discrepancy between model predictions and labels, and distribution information, which has been proven effective in imbalanced classification, is rarely utilized. In contrast, our approach introduces distribution distance optimization on top of the traditional focus on sample-wise prediction error. By concurrently aligning the prediction distribution with the label distribution and minimizing sample-level prediction errors, our method significantly improves accuracy in few-shot regions. This enhancement is achieved without incurring additional computational costs or requiring meticulous hyperparameter tuning. Extensive experiments demonstrate the superiority of our approach in handling critical and informative rare samples in few-shot regions, achieving SOTA results.
@@ 3 Method
Figure 2: The presence of imbalanced data distributions introduces a noticeable distribution discrepancy between the model’s predictions and labels. Dist Loss mitigates this imbalance by simultaneously minimizing this discrepancy and sample-wise prediction errors. Initially, KDE is applied to estimate the label distribution and compute the expected frequency of each label within a batch, thereby generating pseudo-labels that incorporate label distribution information. For example, given the labels [1, 3, 4, 6] and their computed expected frequencies [1, 2, 3, 1], the resulting pseudo-labels would be [1, 3, 3, 4, 4, 4, 6], where each label appears according to its expected frequency. Subsequently, the model’s predictions within a batch are sorted to obtain an ordered sequence that captures the prediction distribution. For instance, if the model’s initial predictions are [5, 2, 6, 3, 2, 7, 1], sorting them yields [1, 2, 2, 3, 5, 6, 7], preserving the distributional characteristics of the predictions. Measuring the distance between these pseudo-labels and pseudo-predictions, which both encapsulate distribution information, provides an approximation of the distributional discrepancy. By optimizing both the distribution distance and sample-level prediction errors during training, the model effectively alleviates the adverse effects of imbalanced data, significantly enhancing accuracy, particularly in few-shot regions.
@@ 3.1 Problem setting
Let $\mathcal{D}$ be a training dataset consisting of $N$ samples, denoted as $\mathcal{D}=\{(\mathbf{x}_{(i)},y_{(i)})\}_{i=1}^{N}$, where $\mathbf{x}_{(i)}\in\mathbb{R}^{d}$ represents the input and $y_{(i)}\in\mathbb{R}$ denotes the corresponding label. To facilitate processing, the continuous label space $\mathcal{Y}$ is discretized into $B$ bins of equal width, such that $\mathcal{Y}=\bigcup_{b=1}^{B}[y_{b},y_{b+1})$, where $y_{b}$ is the lower bound of bin $b$, and $y_{1}<y_{2}<\cdots<y_{B}$. In subsequent discussions, for convenience, the lower bound $y_{b}$ of the bin $[y_{b},y_{b+1})$ will represent any label value $y_{(i)}$ that falls within that bin. The set of bins is denoted as =$\mathcal{B}=\{1,2,\dots,B\}$. In practical scenarios, the width of each bin, denoted $\Delta_{y}$, indicates the minimum resolution of interest when processing the label space. For example, in age estimation, one might set the bin width to 1, resulting in $\Delta_{y}=y_{b+1}-y_{b}=1,\quad\forall b\in\mathcal{B}$. Additionally, we define the probability of observing a label $y_{i}$ as $p_{i}$, and the probability distribution of labels can be estimated using KDE, which is a non-parametric statistical method used to estimate the probability density function of a random variable without assuming a specific distribution form. Notably, the discretization of the label space serves to facilitate the use of label distribution information while preserving the regression nature of the problem, rather than transforming it into a classification task. Since continuous probability density functions are difficult to process directly, we discretize them into bins for practical computation and estimation. This discretization applies to the entire label space, which includes all possible values, not just those occurring in the training dataset.
@@ 3.2 Dist Loss
One of the optimization objectives of Dist Loss is to minimize the distance between the prediction and label distributions in regression tasks. The core challenge lies in measuring the distance between these two distributions in a differentiable manner. Traditional metrics for measuring distribution distance, such as Kullback-Leibler divergence and Jensen-Shannon divergence, cannot be directly implemented in a differentiable form for regression tasks. Therefore, we have devised an alternative approach in the implementation of Dist Loss to realize a differentiable distribution distance measurement in regression scenarios. Specifically, we approximate the distance between the label and prediction distributions by sampling from these distributions and quantifying the differences between the sampled values to estimate the distance.
@@ 3.2.1 Calculation of Dist Loss
To generate pseudo-labels that contain label distribution information, we first randomly sample $M$ points from the label distribution, where M corresponds to the batch size during model training. The expected frequencies of the label $y_{i}$ can be estimated by multiplying the number of sampling points $M$ by the probability of that label $p_{i}$. Based on this, we construct a sequence $\mathcal{N}_{L}=(n_{1},n_{2},\cdots,n_{B})$ to represent these expected frequencies, where $n_{i}=M\cdot p_{i}$. Each element in the obtained $\mathcal{N}_{L}$ represents the expected frequencies of the corresponding label. Since these frequencies may be fractional, we need to convert them to integers while ensuring that the sum after conversion still equals $M$. Here, we denote the converted integer sequence by $\mathcal{N}_{L^{\prime}}=(n^{\prime}_{1},n^{\prime}_{2},\cdots,n^{\prime}_{B})$. To acquire $\mathcal{N}_{L^{\prime}}$, we first take the floor of each element in $\mathcal{N}_{L}$ to obtain the sequence $\mathcal{N}_{L_{f}}=(\lfloor n_{1}\rfloor,\lfloor n_{2}\rfloor,\cdots,\lfloor n_{B}\rfloor)$. Then we calculate the difference $a$, which represents the difference between the sum of the original expected frequencies ($M$) and the sum after applying the floor function, following $a=M-\sum_{i=1}^{B}\lfloor n_{i}\rfloor$. Using the difference $a$, we construct an auxiliary sequence $\mathcal{A}$, which determines how to distribute the difference to the elements of $\mathcal{N}_{L_{f}}$ to ensure the sum is $M$:
@@
Each $n^{\prime}_{i}$ is determined by adding $a_{i}$ to the corresponding element in $\mathcal{N}_{L_{f}}$, where $n^{\prime}_{i}=\lfloor n_{i}\rfloor+a_{i}$, and i $\in\mathcal{B}$. Finally, we generate the corresponding pseudo-labels $\mathcal{S}_{L}$ based on the expected frequencies, here each element $S_{L_{j}}$ is represented as:
$S_{L_{j}}=y_{\,\arg\min_{i\in\mathcal{B}}\left(\sum_{k=1}^{i}n_{k}\geq j\right)},\quad\text{for }j=1,\dots,M,$
where $j=1,\dots,M$ indexes the position in the sequence. To illustrate with a concrete example, consider a label sequence $(y_{1},y_{2},y_{3})=(4,5,6)$ and an obtained frequency sequence $\mathcal{N}_{L^{\prime}}=(1,2,3)$. In this case, the generated pseudo-labels $S_{L}$ would be $(y_{1},y_{2},y_{2},y_{3},y_{3},y_{3})$, corresponding to the sequence $(4,5,5,6,6,6)$.
Similarly, we can perform $M$-point sampling on the prediction distribution and apply the same procedure to obtain the pseudo-predictions $\mathcal{S}_{P}$, which capture the characteristics of the prediction distribution. However, in practice, this process can be simplified by sorting the model predictions within a batch. It is worth noting that although pseudo-labels can also be generated by sorting the labels within a batch, we choose not to adopt this approach here. Instead, we aim to better leverage the full label distribution, as the batch serves merely as a sample. To explain why sorted sequences capture distribution information, the pseudo-labels $\mathcal{S}_{L}$ and pseudo-predictions $\mathcal{S}_{P}$ can be viewed as one-dimensional representations of the label and prediction distributions, respectively (Figure 3).
By measuring the distance between the pseudo-predictions and pseudo-labels, we can approximate the distance between their respective distributions. Let $L(\cdot)$ be a function that measures the distance between two sequences; then, the distribution distance can be expressed as $L(\mathcal{S}_{P},\mathcal{S}_{L})$. Furthermore, using the function $L(\cdot)$, we can simultaneously evaluate the sample-wise prediction errors, which measure the discrepancy between individual predicted values and their corresponding ground truth labels. Specifically, let $Y_{\text{batch}}$ and $\hat{Y}_{\text{batch}}$ denote the sets of ground truth labels and model predictions in a batch, respectively. The sample-wise prediction error can then be formulated as $L(Y_{\text{batch}},\hat{Y}_{\text{batch}})$. Ultimately, by jointly optimizing both the distributional distance and the sample-level prediction errors during training, we can mitigate the issue of distribution mismatch without compromising overall accuracy, thus addressing the challenges posed by imbalanced data distributions.
Figure 3: To illustrate the core concept behind Dist Loss, the figure simplifies its computation by assuming that the batch size equals the total number of training samples.